Solve the separable equation using linear substitution: y,=4/(x+y)2

1
Expert's answer
2021-11-29T11:37:21-0500

Let us solve the separable equation "y'=\frac{4}{(x+y)^2}" using the linear substitution "z=x+y." It follows that "z'=1+y'," and hence we get the equation "z'-1=\frac{4}{z^2}." The last equation is equivalent to "\frac{dz}{dx}=\frac{4}{z^2}+1," and hence to "dx=\frac{z^2dz}{z^2+4}." It follows that

"\int dx=\int\frac{z^2dz}{z^2+4}
=\int\frac{(z^2+4-4)dz}{z^2+4}
=\int dz-4\int\frac{dz}{z^2+4}=z-2\arctan\frac{z}2+C."

Therefore, "x=z-2\arctan\frac{z}2+C."

We conclude that the general solution of the differential equation "y'=\frac{4}{(x+y)^2}" is

"x=x+y-2\arctan\frac{x+y}2+C" or "y=2\arctan\frac{x+y}2+C."